# VRoPE: Rotary Position Embedding for Video Large Language Models

### v1

@@ 5.1 Experimental Setup
We apply our proposed VRoPE to Video-LLM architectures with three widely used LLM backbones: Vicuna-7B, Qwen2-1.5B, and Qwen2-7B, the resulting models are denoted as Video-Vicuna-7B, Video-Qwen2-1.5B, and Video-Qwen2-7B. For the vision encoder, we leverage Eva-CLIP Sun et al. (2023), and connect the Vision Encoder to the LLM using a Multi-Layer Perceptron (MLP) Tolstikhin et al. (2021). For video input, the frames are tokenized using a $2\times 2$ pooling kernel with a stride of 2, i.e., each frame has 64 tokens as input. Training follows a two-stage paradigm: in the pre-training stage, only the MLP connector is trained, while in the instruction-tuning stage, both the MLP and LLM backbones are fine-tuned, with the Vision Encoder frozen throughout. During pre-training, we use a batch size of 256 and a learning rate of 1e-3, while for instruction-tuning, we reduce the batch size to 128 and set the learning rate to 2e-5. A warm-up ratio of 0.03 is used, followed by cosine learning rate decay after the linear warm-up phase. The training was conducted on 8 Nvidia A800 GPUs.
@@
For Vicuna-7B, we pre-train the model on the LLaVA-558K dataset Liu et al. (2024a) with WebVid samples Bain et al. (2021) and fine-tune it on the LLaVA-mix665K Liu et al. (2024a) dataset augmented with VideoChatGPT data Maaz et al. (2023). For the Qwen2 LLM series, we pre-train the models on a randomly sampled 1M caption dataset, which includes LLaVA-558K, WebVid, DenseFusion-1M Li et al. (2024c), VALOR Liu et al. (2024b), and CC3M Changpinyo et al. (2021). The models are then fine-tuned on a combination of LLaVA-mix665K, VideoChatGPT, OneVision Li et al. (2024a), and LLaVA-Video-178K Zhang et al. (2024). We use a $224\times 224$ resolution for both image and video inputs.
@@ 5.2 Main Results
We evaluate the performance of RoPE, RoPE-3D, and our proposed VRoPE across six video understanding benchmarks, with the number of input frames fixed to 16. As shown in Table 1, VRoPE consistently outperforms both RoPE and RoPE-3D, achieving the highest average scores across all tasks and backbones.
For instance, in the Video-Vicuna-7B row, VRoPE achieves an average score of 44.48, surpassing RoPE by 1.13 points. Similarly, when evaluated with larger backbones like Qwen2-1.5B and Qwen2-7B, VRoPE demonstrates consistent improvements across all benchmarks. Notably, it outperforms RoPE and RoPE-3D by significant margins on tasks such as Video-MME (a 3.4-point increase for Qwen2-1.5B) and MLVU (a 2.94-point increase for Qwen2-7B).
These results highlight the superior adaptability of VRoPE across different LLM types and parameter sizes. Importantly, VRoPE introduces no new learnable parameters and does not increase computational complexity, making it a cost-free performance enhancement for Video-LLMs.
5.3 Results on Long Video Retrieval
@@
Table 4: Ablation study on VRoPE components. We evaluate the impact of Cross-Modal Continuity Rotation (Continuity) and Symmetric Bias Mitigation (Symmetric). The model achieves the best performance when both components are applied together.
@@
As shown in Figure 5, the retrieval accuracy of RoPE drops significantly when the number of input frames exceeds 832. While RoPE-3D demonstrates better frame extrapolation capabilities compared to RoPE, VRoPE outperforms both approaches by a considerable margin. The quantitative results, presented in Table 3, further evidence this finding. Specifically, VRoPE achieves an accuracy that is 32.19 points higher than RoPE and 14.22 points higher than RoPE-3D when the number of input frames increases to 1024-1216. Notably, these results are obtained even though the input frame count in this range is dozens of times greater than the maximum number seen during training. This demonstrates the exceptional extrapolation ability of VRoPE.
5.4 Ablation Studies
Comparison of RoPE Variants.
We first compare RoPE-2D Agrawal et al. (2024) and RoPE-3D Wang et al. (2024) with the baseline RoPE Su et al. (2024) method. RoPE-2D encodes only the spatial coordinates $(w,h)$ of each frame, arranged sequentially. While it resolves the cross-modal positional discontinuity of RoPE-3D, it still suffers from positional bias. Both RoPE-2D and RoPE-3D show improvements over RoPE, demonstrating the benefits of spatiotemporal structure modeling.
Next, we evaluate two additional variants, RoPE-Share and RoPE-Compact, to further ablate the impact of S.S.M. and P.U. RoPE-Share uses identical positional embeddings within each frame, arranged sequentially. While it resolves positional bias and ensures continuity, it neglects the spatial structure of the frames, leading to a performance drop compared to RoPE. RoPE-Compact is a extention of RoPE-3D that addresses positional discontinuity by encoding subsequent text tokens with $(W+1,H+1,T+1)$, but it deviates from text compatibility requirements, which slightly limits its performance. In contrast, our proposed method (VRoPE) incorporates all three properties, achieving a 1.99-point improvement over the RoPE baseline, surpassing all other variants.
@@ 5.4 Ablation Studies
We conduct ablation experiments to evaluate the individual contributions of the Cross-Modal Continuity Rotation and Symmetric Bias Mitigation components. The results, presented in Table 4, reveal that when applied separately, each method produces mixed effects. Specifically, Cross-Modal Continuity Rotation improves performance on Video-MME, indicating its effectiveness in enhancing smooth translation for general video understanding. Symmetric Bias Mitigation shows a significant gain on LongVideoBench, indicating its effectiveness in reducing bias in long video tasks. When combined in VRoPE, the two components work synergistically, resulting in more consistent performance.
@@ 7 Limitations
While VRoPE demonstrates strong performance, there are some limitations. Due to computational resource constraints, our experiments were limited to models with 1.5B and 7B parameters. Larger-scale models could potentially yield further performance gains. Additionally, although VRoPE is adaptable across different dimensions, its extension to other modalities (e.g., audio, 3D point clouds, Electroencephalography (EEG)) and higher-dimensional data (e.g., 4D spatiotemporal or medical imaging data) remains an area for future research and validation.

### v2

@@ 5.1 Experimental Setup
We apply our proposed VRoPE to Video-LLM architectures with three widely used LLM backbones: Vicuna-7B, Qwen2-1.5B, and Qwen2-7B, the resulting models are denoted as Video-Vicuna-7B, Video-Qwen2-1.5B, and Video-Qwen2-7B. For the vision encoder, we leverage Eva-CLIP Sun et al. (2023), and connect the Vision Encoder to the LLM using a Multi-Layer Perceptron (MLP) Tolstikhin et al. (2021). We use a $224\times 224$ resolution for both image and video inputs. For video input, the number of input frames is 16 and the frames are tokenized using a $2\times 2$ pooling kernel with a stride of 2, i.e., each frame has 64 tokens as input. Training follows a two-stage paradigm: in the pre-training stage, only the MLP connector is trained, while in the instruction-tuning stage, both the MLP and LLM backbones are fine-tuned, with the Vision Encoder frozen throughout. During pre-training, we use a batch size of 256 and a learning rate of 1e-3, while for instruction-tuning, we reduce the batch size to 128 and set the learning rate to 2e-5. A warm-up ratio of 0.03 is used, followed by cosine learning rate decay after the linear warm-up phase. The training was conducted on 8 Nvidia A800 GPUs.
@@
For Vicuna-7B, we pre-train the model on the LLaVA-558K dataset Liu et al. (2024a) with WebVid samples Bain et al. (2021) and fine-tune it on the LLaVA-mix665K Liu et al. (2024a) dataset augmented with VideoChatGPT data Maaz et al. (2023). For the Qwen2 LLM series, we pre-train the models on a randomly sampled 1M caption dataset, which includes LLaVA-558K, WebVid, DenseFusion-1M Li et al. (2024c), VALOR Liu et al. (2024b), and CC3M Changpinyo et al. (2021). The models are then fine-tuned on a combination of LLaVA-mix665K, VideoChatGPT, and LLaVA-Video-178K Zhang et al. (2024).
@@
5.2 Main Results
We evaluate the performance of RoPE, RoPE-3D, and our proposed VRoPE across six video understanding benchmarks. As shown in Table 2, VRoPE consistently outperforms both RoPE and RoPE-3D, achieving the highest average scores across all tasks and backbones.
For instance, in the Video-Vicuna-7B row, VRoPE achieves an average score of 44.48, surpassing RoPE by 1.13 points. Similarly, when evaluated with Qwen2-1.5B and Qwen2-7B, VRoPE demonstrates consistent improvements across all benchmarks. Notably, it outperforms RoPE and RoPE-3D by significant margins on tasks such as Video-MME (a 3.4-point increase for Qwen2-1.5B) and MLVU (a 2.94-point increase for Qwen2-7B).
These results highlight the superior adaptability of VRoPE across different LLM types and parameter sizes. Importantly, VRoPE introduces no new learnable parameters and does not increase computational complexity, making it a cost-free performance enhancement for Video-LLMs. More results and visualization examples can be found in Appendix B and Appendix C.
@@ 5.2 Main Results
5.3 Results on Long Video Retrieval
As shown in Figure 4, the retrieval accuracy of RoPE drops significantly when the number of input frames exceeds 832, while VRoPE outperforms other approaches by a considerable margin. The quantitative results, presented in Table 4, further evidence this finding. Specifically, VRoPE achieves an accuracy that is 32.19 points higher than RoPE and 14.22 points higher than RoPE-3D when the number of input frames increases to 1024-1216. Notably, these results are obtained even though the input frame count in this range is dozens of times greater than the maximum number seen during training. This demonstrates the exceptional extrapolation ability of VRoPE. Moreover, RoPE-3D underperforms the RoPE baseline for inputs of 256-512, 512-768, and 768-1024 frames, which further proves that the cross-modal positional discontinuity affects the model’s ability to understand videos of different lengths.
5.4 Ablation Studies
Comparison of RoPE Variants.
We first compare RoPE-2D Agrawal et al. (2024) and RoPE-3D Wang et al. (2024) with the baseline RoPE Su et al. (2024) method. RoPE-2D encodes only the spatial coordinates $(w,h)$ of each frame. While it resolves the cross-modal positional discontinuity, it still suffers from positional bias. Both RoPE-2D and RoPE-3D show improvements over RoPE, demonstrating the benefits of spatiotemporal structure modeling.
Next, we evaluate two additional variants, RoPE-Share and RoPE-Compact, to further ablate the impact of S.S.M. and P.U. RoPE-Share uses identical positional embeddings within each frame, arranged sequentially. While it resolves positional bias and ensures continuity, it neglects the spatial structure of the frames, leading to a performance drop compared to RoPE. RoPE-Compact is an extention of RoPE-3D that addresses positional discontinuity by encoding subsequent text tokens with $(W+1,H+1,T+1)^{T}$, but it deviates from text compatibility requirements, which slightly limits its performance. In contrast, our proposed method (VRoPE) incorporates all three properties, achieving a 1.99-point improvement over the RoPE baseline, surpassing all other variants. More detailed illustration of RoPE-Share and RoPE-Compact can be found in Appendix D.
Table 5: Ablation study on VRoPE components. We evaluate the impact of Symmetric Bias Mitigation (Symmetric) and Temporal Centered Arrangement (Continuity). The model achieves the best performance when both components are applied together.
@@ 5.4 Ablation Studies
We conduct ablation experiments to evaluate the individual contributions of the Symmetric Bias Mitigation and Temporal Centered Arrangement components. The results, presented in Table 5, reveal that when applied separately, each method produces mixed effects. Specifically, Temporal Centered Arrangement improves performance on Video-MME, indicating its effectiveness in enhancing smooth translation for general video understanding. Symmetric Bias Mitigation shows a significant gain on LongVideoBench, indicating its effectiveness in reducing bias in long video tasks. When combined in VRoPE, the two components work synergistically, resulting in more consistent performance.
@@ 7 Limitations
While VRoPE demonstrates strong performance, there are some limitations. Due to computational resource constraints, our experiments were limited to models with 1.5B, 7B and 8B (shown in Appendix B) parameters. Larger-scale models could potentially yield further performance gains. Additionally, although VRoPE is adaptable across different dimensions, its extension to other modalities (e.g., audio, 3D point clouds, Electroencephalography (EEG)) and higher-dimensional data (e.g., 4D spatiotemporal or medical imaging data) remains an area for future research and validation.
Appendix A Discussion
Dimensional Adaptability.
A key advantage of VRoPE is its ability to degenerate into lower-dimensional embeddings without altering its fundamental structure. Unlike methods that assign separate feature channels for each coordinate, VRoPE employs linear combinations of the original coordinates, allowing any dimension set to 1 to seamlessly adapt into lower-dimension form. For instance, when $H=1$, the encoded positions simplify to $(w,w,-w,-w)$, effectively reducing to a 1D form—unlike previous methods that rely on separate encodings, such as $(w,0)$. This property is particularly beneficial for adapting pre-trained model’s positional encodings from images (2D) or videos (3D) to data of varying dimensions without disrupting the original encoding scheme. Consequently, models can transfer more effectively across modalities while preserving consistent positional behavior.
Appendix B More Results
B.1 Results on EventBench
Table 6: Performance comparison of RoPE variants on event-based EventBench Du et al. (2024).
Method EventBench
Video-Vicuna-7B
    w/ RoPE 38.97
    w/ RoPE-3D 39.33 ($\uparrow$0.36)
    w/ VRoPE 40.38 ($\uparrow$1.41)
Video-Qwen2-1.5B
    w/ RoPE 53.31
    w/ RoPE-3D 52.76 ($\downarrow$0.55)
    w/ VRoPE 54.23 ($\uparrow$0.92)
Video-Qwen2-7B
    w/ RoPE 59.25
    w/ RoPE-3D 58.61 ($\downarrow$0.64)
    w/ VRoPE 60.35 ($\uparrow$1.1)
The benchmark evaluated in Section 5.2 already encompasses comprehensive capabilities required for video understanding tasks. To further validate temporal reasoning performance, we conduct additional evaluations focusing on event-based tasks involving complex temporal dependencies. As shown in Table 6, our VRoPE demonstrates consistent improvements across all models compared to existing methods. These results confirm that our approach maintains superior comprehension capabilities when processing videos containing intricate event sequences.
B.2 Results on Video-MME with varying lengths
Table 7: Detailed performance comparison of RoPE variants on Video-MME Fu et al. (2024).
Method Short Medium Long
Video-Vicuna-7B
    w/ RoPE 46.4 38.0 31.0
    w/ RoPE-3D 46.0 ($\downarrow$0.4) 37.5 ($\downarrow$0.5) 30.6 ($\downarrow$0.4)
    w/ VRoPE 46.4 (-) 38.3 ($\uparrow$0.3) 31.8 ($\uparrow$0.8)
Video-Qwen2-1.5B
    w/ RoPE 47.4 37.6 32.2
    w/ RoPE-3D 47.1 ($\downarrow$0.3) 37.0 ($\downarrow$0.6) 33.8 ($\uparrow$1.6)
    w/ VRoPE 50.1 ($\uparrow$2.7) 39.3 ($\uparrow$1.7) 37.8 ($\uparrow$5.6)
Video-Qwen2-7B
    w/ RoPE 60.2 47.6 42.5
    w/ RoPE-3D 60.0 ($\downarrow$0.2) 46.7 ($\downarrow$0.9) 41.7 ($\downarrow$0.8)
    w/ VRoPE 60.4 ($\uparrow$0.2) 47.6 (-) 43.9 ($\uparrow$1.4)
In this section, we analyze the performance of RoPE, RoPE-3D, and our VRoPE across varying input video lengths on the Video-MME dataset, as summarized in Table 7. The results indicate that VRoPE demonstrates marked superiority in processing long-form videos, while also achieving moderate advantages for medium and short videos, maintaining comparable performance to baselines at minimum. This further validates the effectiveness of our approach in enhancing model comprehension capabilities across varying video durations. The consistent improvements underscore our method’s robustness in understanding tasks under various video context lengths.
B.3 Results under Challenging Conditions
Table 8: Results on Video-MME Du et al. (2024) under lower frame rates (8 frames).
Method Acc.
Video-Qwen2-1.5B
    w/ RoPE 38.9
    w/ RoPE-3D 37.2 ($\downarrow$1.7)
    w/ VRoPE 40.9 ($\uparrow$2.0)
Figure 5: Attention weight visualization of RoPE, RoPE-3D, and VRoPE. The visualization reveals that VRoPE exhibits stronger attention activation within critical frames (highlighted by red boxes), demonstrating its accurate focus on pivotal spatiotemporal regions. In contrast, RoPE and RoPE-3D display attenuated attention responses in these corresponding areas, indicating insufficient awareness of key events. This attention misalignment consequently leads to erroneous predictions, as evidenced by their incorrect interpretations of the visual content.
In this section, we evaluate the performance of RoPE, RoPE-3D, and VRoPE on Video-MME under low frame-rate inputs (8 frames), as reported in Table 8. Notably, VRoPE maintains enhanced performance even in these challenging sparse-sampling scenarios, empirically confirming the robustness of our approach. This empirical evidence highlights our method’s capability to preserve spatiotemporal coherence under severe input degradation.
Table 9: Performance comparison of RoPE variants on larger models and datasets. Results across tasks, including general video understanding (Video-MME), video temporal understanding (MVBench, TempCompass), and long video understanding (MLVU, LongVideoBench, EgoSchema).
(w/o subs) @M-Avg @Val @Multi-Choice @Test
Video-Qwen3-8B
    w/ RoPE 61.00 64.96 59.68 60.81 68.67 56.41 61.92
    w/ RoPE-3D 61.44 ($\uparrow$0.44) 64.50 ($\downarrow$0.46) 59.34 ($\downarrow$0.34) 61.00 ($\uparrow$0.19) 69.11 ($\uparrow$0.44) 56.03 ($\downarrow$0.38) 61.90 ($\downarrow$0.02)
    w/ VRoPE 62.56 ($\uparrow$1.56) 65.36 ($\uparrow$0.40) 59.23 ($\downarrow$0.45) 61.48 ($\uparrow$0.67) 68.67 (-) 57.07 ($\uparrow$0.66) 62.40 ($\uparrow$0.48)
B.4 Results of Larger Models and Datasets
In this section, we validate the superiority of our approach through scaled-up model architectures and expanded training datasets. Specifically, we conduct experiments using SigLIP-2 Tschannen et al. (2025) and Qwen3-8B Yang et al. (2025) as backbone architectures. We expand the number of input frames to 32 and the resolution is set to $384\times 384$. During the pre-training stage, we utilize LLaVA-558K Liu et al. (2024a) combined with 500K randomly sampled video-text pairs from OpenVid-1M Nan et al. (2024). For instruction tuning, we integrate LLaVA-NeXT-790K Li et al. (2024a), LLaVA-Video-178K Zhang et al. (2024), and the full OpenVid-1M dataset. This configuration results in approximately 1 million samples for pre-training and 3 million samples for instruction tuning. As demonstrated in Table 9, VRoPE maintains performance advantages even under these enhanced baseline conditions (larger models, expanded datasets, and stronger baselines). These results further substantiate the generalizability and robustness of our method across diverse architectural scales and data regimes.
Appendix C Visualization Analysis
In Section 3.4, we analyze the positional attention bias and cross-modal positional discontinuity inherent to RoPE and RoPE-3D. To further substantiate these observations, we provide concrete attention visualization examples in this section. As illustrated in Figure 5, for an input video sequence, our VRoPE effectively focuses on the video frames most relevant to the query (the red door and the helicopter), whereas RoPE and RoPE-3D exhibit insufficient attention to critical frames. This deficiency leads to localization errors and subsequent incorrect responses – for instance, misidentifying the opening of a black door as the opening of a red door in this example. The comparative visualization demonstrates our method’s enhanced capability in spatiotemporal feature localization and event understanding.
Appendix D Detailed Illustration of Other RoPE Variants
RoPE-Share.
RoPE-Share is a 1D positional encoding where all spatial tokens within a video frame share the same positional ID, i.e., the positional IDs of all frame tokens in the $t$th frame are $n+t$. Text tokens follow the original encoding: $n+T+1,n+T+2,...$. While this design eliminates spatial attention bias and ensures cross-modal continuity, it fails to model spatial positional relationships within frames, leading to suboptimal performance (as is shown in Section 5.4).
RoPE-Compact.
RoPE-Compact is a variant of RoPE-3D. The key difference lies in handling cross-modal boundaries: (1) RoPE-3D assigns the next text token a positional ID of $(\max(W,H,T),\max(W,H,T),\max(W,H,T))^{T}$. For example, if $T>W,H$, the last video token is $(W,H,T)^{T}$, and the next text token becomes $(T,T,T)^{T}$, causing discontinuity in the $w$ and $h$ dimensions (as shown in Section 5.3). (2) To address the above issue, RoPE-Compact increments each dimension by 1, and uses it as the positional ID for the next text token: $(W+1,H+1,T+1)^{T}$. While this resolves cross-modal discontinuity, it disrupts the pre-trained RoPE’s positional frequency patterns of text, degrading performance.
Appendix E License Statement
The scientific artifacts used in this work are all publicly available and this work only uses them for research purposes, thus not violating any of the artifacts’ licenses. The new models released in this work is also licensed for research purposes only, prohibiting any other misuse.